OpenStudy (anonymous):

he pressure of a gas p(v) varies inversely with the volume of the gas v. The pressure of a gas measures 15 kg/cm2 when its volume is 300 cm3. Which equation can be used to find the pressure of the gas when the volume is changed?

OpenStudy (anonymous):

p(v) = 20v p(v) = 20/v p(v) = 4500v p(v)= 4500 / v

OpenStudy (anonymous):